OpenStudy (anonymous):

Find the width of a rectangle with a perimeter of 90 and a length of 15

myininaya (myininaya):

perimeter=P=2L+2W=90 length=L=15 so we have 2L+2W=90 and L=15 2(15)+2W=90 solve for W

OpenStudy (anonymous):

2w = 90 - 2*15 = 60 w = 30 width is greater than length maybe they shpould be reversed

myininaya (myininaya):

15+W=90/2 (divided both sides by 2) 15+W=45 W=45-15=30

OpenStudy (radar):

x=width 2x+30=90 2x=60 x=30 width = 30